OpenStudy (blu_seaa):

Evaluate. 25 23 A) 1 B) 28 C) 4 D) 8

OpenStudy (unklerhaukus):

?

OpenStudy (blu_seaa):

I did that by accident.

OpenStudy (blu_seaa):

Sorry.

OpenStudy (unklerhaukus):

is that 2^5 / 2^3 ?

OpenStudy (blu_seaa):

Yes.

OpenStudy (blu_seaa):

B is an exponent as well.

OpenStudy (unklerhaukus):

(2 x 2 x 2 x 2 x 2) / (2 x 2 x 2)

OpenStudy (unklerhaukus):

cancel what you can , what is left?

OpenStudy (blu_seaa):

Uh, 4046?

OpenStudy (blu_seaa):

I just did what you told me to do. (2 x 2 x 2 x 2 x 2) / (2 x 2 x 2)

OpenStudy (unklerhaukus):

\[\frac{2\times2\times2\times2\times2}{2\times2\times2}=\frac{2\times2\times\cancel{2\times2\times2}}{\cancel{2\times2\times2}}\]

OpenStudy (blu_seaa):

So it's C?

OpenStudy (unklerhaukus):

the other way is to use the index law \[\frac{x^m}{x^n}=x^{m-n}\]

OpenStudy (skullpatrol):

yes, C is the answer

OpenStudy (blu_seaa):

Thank you.

OpenStudy (unklerhaukus):

yep

OpenStudy (skullpatrol):

Good job!